### v2

# Automorphism groups and Distinguishing Colorings of Central and Middle Graphs

Abstract.
Let $G$ be a simple, finite, connected, and undirected graph. The middle graph $M(G)$ of $G$ is obtained from the subdivision graph $S(G)$ after joining pairs of subdivided vertices that lie on adjacent edges of $G$ and the central graph $C(G)$ of $G$ is obtained from $S(G)$ after joining all non-adjacent vertices of $G$. We show that if the order of $G$ is at least $4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$(as abstract groups) and apply these results to obtain new upper bounds of the distinguishing number and the distinguishing index of $C(G)$ and $M(G)$ inspired by an algorithm by Kalinowski, Pilśniak, and Woźniak from 2016.
Key words and phrases:
Automorphism group, distinguishing coloring, central graph, middle graph, subdivision graph
2020 Mathematics Subject Classification:
Primary 05C15; Secondary 05C25, 05C76.
Corresponding author. In this revised version, all results except Propositions 3.2 and 3.3 were obtained and written by the first author. The ideas underlying Propositions 3.2 and 3.3 originate from an earlier version of the manuscript by all authors, where they appeared in the form of diagrams. The propositions are formalized and proved here by the first author.
1. Introduction
All graphs considered in this paper are finite, simple, connected, and undirected. Let $G=(V(G),E(G))$ be a connected graph with the vertex set $V(G)$ and the edge set $E(G)$. Let $\Delta(G)$ represent the maximum degree of $G$. Let $C_{n}$ be a cycle graph of $n$ vertices, $P_{n}$ be a path of $n$ vertices, $K_{n}$ be a complete graph of $n$ vertices, and $K_{n,m}$ be a complete bipartite graph with bipartitions of sizes $n$ and $m$. A claw is the graph $K_{1,3}$. The degree of a vertex $v$ in $G$, denoted by $deg_{G}(v)$, is the number of edges that emerge from $v$. We say $G$ is a regular graph if all vertices of $G$ have the same degree. The degree of a regular graph $G$, denoted by $d(G)$, is the same as the maximum degree $\Delta(G)$ and the minimum degree $\delta(G)$. Let $N_{G}(v)=\{u\in V(G):uv\in E(G)\}$ for any $v\in V(G)$. The middle graph $M(G)$ and the central graph $C(G)$, introduced by Hamada and Yoshimura [8] and Vernold [18], respectively, are classes of graphs derived from subdivision graphs. These graph constructions have been extensively studied; see, for example, [4, 5, 6, 7, 11, 12, 13, 15].
Definition 1.1.
The subdivision graph $S(G)$ of $G$ is obtained from $G$ by putting a new vertex in the middle of every edge of $G$. For each edge $xy\in E(G)$, introduce a new vertex $w_{xy}$ (called the subdivision vertex corresponding to the edge $xy$) and replace the edge $xy$ by the two edges $xw_{xy}$ and $w_{xy}y$. Thus, $V(S(G))=V(G)\cup\{w_{xy}:xy\in E(G)\},$ and $E(S(G))=\{xw_{xy},\,w_{xy}y:xy\in E(G)\}.$ The middle graph $M(G)$ of $G$ is the graph whose vertex set is $V(G)\cup E(G)$, and two vertices $x,y$ in the vertex set of $M(G)$ are adjacent in $M(G)$ in the case where one of the following holds:
  1. (1)
    $x,y\in E(G)$ and $x,y$ are adjacent in $G$;
  2. (2)
    $x\in V(G),y\in E(G)$, and $x,y$ are incident in $G$.
The central graph $C(G)$ of $G$ with order $n$ and size $m$ is a graph of order $n+m$ and size $n\choose{2}$ + $m$ that is obtained from $S(G)$ by adding the edges in $E(\overline{G})$, where $\overline{G}$ is the complement graph of $G$.
Definition 1.2.
Let $Aut(G)$ be the automorphism group of $G$, $id_{G}$ be the trivial automorphism of $G$, and $f$ be a vertex coloring or an edge coloring or a total coloring (that is, a coloring of its vertices and edges) of $G$. An automorphism $\phi$ of $G$ preserves $f$ if the following holds:
  1. (1)
    $f(x)=f(\phi(x))$ for all $x\in V(G)$ if $f$ is a vertex coloring,
  2. (2)
    $f(e)=f(\phi(e))$ for all $e\in E(G)$ if $f$ is an edge coloring,
The notions of distinguishing number and distinguishing index were introduced by Albertson and Collins [1] and Kalinowski and Pilsniak [10], respectively.
Definition 1.3.
The distinguishing number (index) of a graph $G$, denoted by $D(G)$ ($D^{\prime}(G)$), is the least integer $d$ such that $G$ admits a vertex coloring (edge coloring) with $d$ colors that is preserved only by $id_{G}$.
1.1. Motivation
Alikhani and Soltani [3, Theorems 3.8 and 4.3] studied the distinguishing number and distinguishing index of subdivision graphs and proved that if $G$ is a connected graph of order $n\geq 3$, then
$D(S(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil\quad\text{and}\quad D^{\prime}(S(G))\leq\min\{r:r^{2}-r\geq D^{\prime}(G)\},$
where $\Delta(G)$ denotes the maximum degree of $G$. Kalinowski–Pilśniak–Woźniak [9] investigated the total distinguishing chromatic number of graphs. Inspired by the results of Alikhani and Soltani [3] and the results of Kalinowski–Pilśniak–Woźniak [9], we investigate the distinguishing parameters of middle and central graphs.
1.2. Main Results
Let $G$ be a connected graph of order $n$. We prove the following:
  1. (1)
    (Theorem 2.9) If $n\geq 4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$.
  2. (2)
    (Theorem 3.10(1)) If $n\geq 3$, then $D(M(G))\leq\Delta(G)$;
  3. (3)
    (Theorem 3.10(2)) $D^{\prime}(M(G))\leq 3$;
  4. (4)
    (Theorem 3.4) If $n\geq 4$, then $D^{\prime}(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil.$
  5. (5)
    (Theorem 3.8) If $n\geq 4$, then $D(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil.$
2. Automorphism groups
In this section, we sketch some lemmas concerning $Aut(C(G))$ and $Aut(M(G))$ where $M(G)$ is the middle graph of $G$. Applying these lemmas, we will prove Theorems 3.4 and 3.8 and propositions 3.2, 3.3, and 3.6. Recall that for each edge $xy\in E(G)$, $w_{xy}$ denotes the subdivision vertex corresponding to the edge $xy$ in the subdivision graph $S(G)$. In $S(G)$, let $V_{1}=V(G)$ be the set of all vertices of $G$, and let $V_{2}=\{w_{xy}:xy\in E(G)\}$ denote the set of all subdivision vertices.
Lemma 2.1.
If $\psi\in Aut(C(G))$ such that $\psi(u)=u$ for all $u\in V_{1}$, then $\psi=id_{C(G)}$.
Proof.
First, we observe that distinct vertices of $V_{2}$ have distinct neighborhoods in $C(G)$. Indeed, if $w_{xy},w_{uv}\in V_{2}$ with $w_{xy}\neq w_{uv}$, then $\{x,y\}\neq\{u,v\}$ since $G$ is simple, and hence each edge of $G$ has a unique pair of endpoints. Therefore,
$N_{C(G)}(w_{xy})=\{x,y\}\neq\{u,v\}=N_{C(G)}(w_{uv}).$
Now, let $x\in V_{2}$. If $\psi(x)=y$ for some $y\in V_{2}\setminus\{x\}$, then $\psi(N_{C(G)}(x))=N_{C(G)}(y)$ since automorphisms preserve adjacency. Since $N_{C(G)}(x)\subseteq V_{1}$ and $\psi(v)=v$ for every $v\in V_{1}$, it follows that $\psi(N_{C(G)}(x))=N_{C(G)}(x)$. Thus, $N_{C(G)}(x)=N_{C(G)}(y)$, a contradiction. Since $\psi(u)=u$ for all $u\in V_{1}$, we have $\psi(V_{1})=V_{1}$, and hence $\psi(x)\notin V_{1}$. Therefore, $\psi(x)=x$ for all $x\in V_{2}$, and consequently $\psi=id_{C(G)}$. ∎
Lemma 2.2.
Let $G$ be a connected graph of order $n\geq 4$. If $\varphi\in Aut(C(G))$, then $\varphi(V_{1})=V_{1}$ and $\varphi(V_{2})=V_{2}$.
Proof.
In $C(G)$, each vertex of $V_{1}$ has degree $n-1\geq 3$, whereas each vertex of $V_{2}$ has degree $2$. Since automorphisms preserve vertex degrees, that is, $deg_{C(G)}(\varphi(v))=deg_{C(G)}(v)$ for any $v\in V(C(G))$, it follows that $\varphi(V_{1})=V_{1}$ and $\varphi(V_{2})=V_{2}$. ∎
Lemma 2.3.
Let $G$ be a connected graph of order $n\geq 4$. Then $Aut(G)\cong Aut(C(G))$.
Proof.
Let $\alpha\in Aut(G)$. Define the mapping $\pi_{\alpha}:V(C(G))\rightarrow V(C(G))$ as follows:
$\pi_{\alpha}(w)=\begin{cases}\alpha(w)&\text{if }w\in V(G)\\ w_{\alpha(x)\alpha(y)}&\text{if }w\in V(C(G))\backslash V(G),w=w_{xy},x,y\in V(G).\end{cases}$
We can see that $\pi_{\alpha}\in Aut(C(G))$. If $xy\in E(C(G))$ then either $xy\in E(\overline{G})$ or $xy\in E(S(G))$. By the definition of $\pi_{\alpha}$, if $xy\in E(S(G))$ then $\pi_{\alpha}(x)\pi_{\alpha}(y)\in E(S(G))$. Moreover, we have
$xy\in E(\overline{G})$ $\iff xy\notin E(G)$
$\iff\alpha(x)\alpha(y)\notin E(G)$
$\iff\alpha(x)\alpha(y)\in E(\overline{G})$
$\iff\pi_{\alpha}(x)\pi_{\alpha}(y)\in E(\overline{G}).$
Thus, if $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$, then $H\subseteq Aut(C(G))$. We show that $Aut(C(G))\subseteq H$. Let $f\in Aut(C(G))$. By Lemma 2.2, we have $f(V_{1})=V_{1}$ since $G$ is a connected graph of order at least $4$. Thus, if $\alpha=f\restriction V_{1}$, then $\alpha\in Aut(G)$. Let $\pi_{\alpha}$ be the induced automorphism of $\alpha$ as defined above. If we let $l=f^{-1}\circ\pi_{\alpha}$, then $l$ is an automorphism of $C(G)$ such that $l(v)=v$ for every $v\in V_{1}$. By Lemma 2.1, $l=id_{C(G)}$. Consequently, $f=\pi_{\alpha}$. Therefore, $Aut(C(G))\subseteq H$, and so $H=Aut(C(G))$. It is trivial to see that $H\cong Aut(G)$, so we are done. ∎
The following definition is due to Hamada and Yoshimura [8].
Definition 2.4.
[…]
Fact 2.5.
The following holds:
  1. (1)
    (Hamada–Yoshimura; [8]) Let $G$ be any graph. Then $M(G)\cong L(G^{+})$.
  2. (2)
    (Sabidussi; [17]) If $G$ is a connected graph where $G\not\in\{P_{2},Q,L(Q)\}$ (see Fig. 1). Then $Aut(G)\cong Aut(L(G))$ where $L(G)$ is the line graph of $G$.
  3. (3)
    (Mirafzal; [14, Theorem 4]) If $G\not\cong C_{n}$ for any $n\geq 3$, then $Aut(S(G))\cong Aut(G)$.
[…]
Figure 1. Graphs $L(Q),Q,$ and $C_{6}^{+}$.
Lemma 2.6.
If $\psi\in Aut(G^{+})$ such that $\psi(u)=u$ for all $u\in U_{1}$, then $\psi=id_{G^{+}}$.
Proof.
Let $u_{i}\in U_{2}$. Suppose $\psi(u_{i})=u_{j}$ for some $u_{j}\in U_{2}$ with $j\neq i$. Since automorphisms preserve adjacency and $u_{i}$ is adjacent only to $v_{i}\in U_{1}$, it follows that $\psi(v_{i})=v_{j}$. However, by assumption $\psi(v_{i})=v_{i}$ for all $v_{i}\in U_{1}$, which is a contradiction. Thus $\psi(u_{i})=u_{i}$ for all $u_{i}\in U_{2}$, and hence $\psi=id_{G^{+}}$. ∎
Lemma 2.7.
If $G$ is a non-trivial connected graph and $\varphi\in Aut(G^{+})$, then $\varphi(U_{i})=U_{i}$ for $i\in\{1,2\}$.
Proof.
In $G^{+}$, each vertex of $U_{1}$ has degree at least $2$, while each vertex of $U_{2}$ has degree $1$. Since automorphisms preserve vertex degrees, it follows that $\varphi(U_{1})=U_{1}$ and $\varphi(U_{2})=U_{2}$. ∎
Lemma 2.8.
Let $G$ be a connected graph of order $n\geq 3$. Then $Aut(G)\cong Aut(M(G))$.
Proof.
Since $G^{+}\not\in\{Q,L(Q),P_{2}\}$ as $G^{+}$ has $n$ vertices of degree one (see Fig. 1), $Aut(M(G))\cong Aut(L(G^{+}))$ by the Fact 2.5(1), and $Aut(L(G^{+}))\cong Aut(G^{+})$ by the Fact 2.5(2). Let $\alpha\in Aut(G)$. Define the mapping $\pi_{\alpha}:V(G^{+})\rightarrow V(G^{+})$ by the following rule:
$\pi_{\alpha}(w)=\begin{cases}\alpha(v_{i})&\text{if }w=v_{i}\in U_{1}\\ u_{j}&\text{if }w=u_{i}\in U_{2},v_{j}=\alpha(v_{i}).\end{cases}$
Then $\pi_{\alpha}\in Aut(G^{+})$, where $\pi_{\alpha}(U_{i})=U_{i}$ for $i\in\{1,2\}$. The rest follows by applying Lemmas 2.6, 2.7, and the arguments in the proof of Lemma 2.3. ∎
By Lemmas 2.3 and 2.8, together with Fact 2.5(2,3), we obtain the following result.
Theorem 2.9.
If $G$ is a connected graph of order $n\geq 5$ and $G\not\cong C_{n}$, then $Aut(G)\cong Aut(L(G))\cong Aut(S(G))\cong Aut(C(G))\cong Aut(M(G))$.
Proof.
Since $n\geq 5$, we have $G\not\cong P_{2},Q,$ or $L(Q)$. Hence, by Fact 2.5(2), $Aut(G)\cong Aut(L(G)).$ Since $G\not\cong C_{n}$, Fact 2.5(3) implies that $Aut(S(G))\cong Aut(G).$ Moreover, by Lemma 2.3 and Lemma 2.8, we have $Aut(C(G))\cong Aut(G)\text{ and }Aut(M(G))\cong Aut(G).$
3. Distinguishing number and distinguishing index
In [9], Kalinowski, Pilśniak, and Woźniak introduced the total distinguishing number $D^{\prime\prime}(G)$ of a graph $G$ as the least integer $d$ such that $G$ has a total coloring (not necessarily proper) with $d$ colors that is preserved only by $id_{G}$.
Fact 3.1.
Let $G$ be a connected graph of order $n$. The following holds:
  1. (1)
    (Kalinowski, Pilśniak, and Woźniak; [9, Theorem 2.2]) If $n\geq 3$, then $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$.
  2. (2)
    (Alikhani and Soltani; [3, Theorems 3.8 and 4.3]) If $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$.
  3. (3)
    (Kalinowski and Pilśniak; [10, Propositions 5 and 6]) $D^{\prime}(C_{t})=2$ and $D^{\prime}(K_{t})=2$ for $t\geq 6$.
  4. (4)
    (Alikhani and Soltani; [2, Theorem 2.3]) If $G\not\in\{P_{2},L(Q)\}$, then $D(L(G))=D^{\prime}(G)$ (see Fig. 1).
  5. (5)
    (Kalinowski and Pilśniak; [10]) If $n\geq 3$, then $D^{\prime}(G)\leq\Delta(G)$ unless $G$ is isomorphic to $C_{3},C_{4},$ or $C_{5}$.
Proposition 3.2.
For any integer $n\geq 4$, we have $D^{\prime}(C(K_{n}))=2$.
Proof.
By Lemma 2.3, we have $Aut(C(K_{n}))\cong Aut(K_{n})\cong S_{n}$. Thus, $|Aut(C(K_{n}))|=n!>1$ and so $C(K_{n})$ is not asymmetric. Consequently, $D^{\prime}(C(K_{n}))>1$ and it suffices to construct an edge coloring $c:E(C(K_{n}))\to\{1,2\}$ such that the only automorphism of $C(K_{n})$ that preserves $c$ is $id_{C(K_{n})}$. Let $V(K_{n})=\{v_{1},\dots,v_{n}\}$. Recall that for each edge $v_{i}v_{j}$ of $K_{n}$, $w_{v_{i}v_{j}}$ is the subdivision vertex introduced on that edge. Define the edge coloring $c:E(C(K_{n}))\to\{1,2\}$ such that for each $v_{i}\in V_{1}$, the incident edges $v_{i}w_{v_{i}v_{j}}$ are colored $2$ if $j<i$ and $1$ if $j>i$ (see Figure 2).
[…]
Figure 2. A schematic representation of the edge coloring of $C(K_{n})$. Edges $v_{i}w_{v_{i}v_{j}}$ are colored 22 (blue) if $j<i$ and 11 (red) if $j>i$.
Let $\varphi\in\mathrm{Aut}(C(K_{n}))$ be an automorphism that preserves $c$. For each $i\in\{1,...,n\}$, let $d_{2}(v_{i})$ denote the number of edges incident to $v_{i}$ that are colored with $2$. By construction, $d_{2}(v_{i})=i-1\text{ for all }i\in\{1,\dots,n\}.$ Hence, the values $d_{2}(v_{i})$ are distinct. Since $\varphi$ preserves $c$, we have $\varphi(v_{i})=v_{i}$ for all $v_{i}\in V(K_{n})$. Thus, by Lemma 2.1, $\varphi=id_{C(K_{n})}$. ∎
Proposition 3.3.
For any integer $n\geq 4$, we have $D^{\prime}(C(C_{n}))=2$.
Proof.
Since $Aut(C(C_{n}))\cong Aut(C_{n})\cong D_{n}$ by Lemma 2.3 where $D_{n}$ is the Dihedral group, the graph $C(C_{n})$ is not asymmetric, and thus $D^{\prime}(C(C_{n}))>1$. We show $D^{\prime}(C(C_{n}))\leq 2$. Let $V(C(G))=V_{1}\cup V_{2}$, where $V_{1}=V(C_{n})$ and $V_{2}=\{w_{xy}:xy\in E(C_{n})\}$ be the set of subdivision vertices. Since $n\geq 4$, we can select distinct $v_{1},v_{2},v_{3}\in V_{1}$ so that $v_{1}v_{2},v_{2}v_{3}\in E(C_{n})$. Define an edge coloring $c:E(C(C_{n}))\to\{0,1\}$ as follows (see Figure 3):
  1. (1)
    $c(v_{1}w_{v_{1}v_{2}})=1$ and $c(w_{v_{1}v_{2}}v_{2})=1$.
  2. (2)
    $c(v_{3}w_{v_{2}v_{3}})=1$.
  3. (3)
    $c(e)=0$ otherwise.
Let $\phi\in Aut(C(C_{n}))$ be an automorphism that preserves $c$. Let $d_{1}(u)$ denote the number of edges incident to $u$ colored with $1$. We can observe the following:
  • The vertices $w_{v_{1}v_{2}}$ and $w_{v_{2}v_{3}}$ are the unique vertices with $d_{1}(w_{v_{1}v_{2}})=2$ and $d_{1}(w_{v_{2}v_{3}})=1$. So $\phi(w_{v_{1}v_{2}})=w_{v_{1}v_{2}}$ and $\phi(w_{v_{2}v_{3}})=w_{v_{2}v_{3}}$. Since automorphism preserves neighbors, we have $\phi(\{v_{1},v_{2}\})=\{v_{1},v_{2}\}$ and $\phi(\{v_{2},v_{3}\})=\{v_{2},v_{3}\}$.
  • Among the neighbors $v_{2}$ and $v_{3}$ of $w_{v_{2}v_{3}}$, $v_{3}$ is the only vertex with $d_{1}(v_{3})=1$.111note that $d_{1}(v_{2})$ also is $1$, but $v_{2}\in N(w_{v_{1}v_{2}})$ Thus $\phi(v_{3})=v_{3}$. Consequently, $\phi(v_{2})=v_{2}$. Since $\phi(\{v_{1},v_{2}\})=\{v_{1},v_{2}\}$, we have $\phi(v_{1})=v_{1}$.
Since $\phi$ fixes three consecutive vertices $v_{1},v_{2},$ and $v_{3}$ of the cycle $C_{n}$, $\phi$ must fix all vertices of $C_{n}$. By Lemma 2.1, $\phi=id_{C(C_{n})}$. ∎
[…]
Figure 3. A schematic representation of the edge coloring of $C(C_{n})$. The edges $v_{1}w_{v_{1}v_{2}},w_{v_{1}v_{2}}v_{2}$, and $v_{3}w_{v_{2}v_{3}}$ are colored with 11 (red), and the rest of the edges are colored with 0 (black).
Theorem 3.4.
Let $G$ be a connected graph of order $n\geq 4$. Then $D^{\prime}(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$.
Proof.
Since $G$ is connected, we have $\Delta(G)\geq 2$. Hence,
$\left\lceil\sqrt{\Delta(G)}\right\rceil\geq\left\lceil\sqrt{2}\right\rceil=2.$
For $k\geq 0$, let $S_{k}(x)$ denote a sphere of radius $k$ with a center $x$ in $G$, that is, the set of all vertices at distance $k$ from $x$.
Case (i): $G=K_{n}$ or $C_{n}$.
By Propositions 3.2 and 3.3, $D^{\prime}(C(G))=2$. Since $\left\lceil\sqrt{\Delta(G)}\right\rceil\geq 2,$ we obtain $D^{\prime}(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil$.
Case (ii): $G$ contains a cycle but $G\not\in\{C_{n},K_{n}\}$.
Thus, we can choose a vertex $v_{0}$ lying on a cycle such that the sphere $S_{2}(v_{0})$ is non-empty. Consider a BFS-spanning tree $T$ of $G$ rooted at $v_{0}$. For a vertex $v\in T$, we denote
$X_{v,u}=(\{v,w_{vu}\},\{w_{vu},u\})$ if $w_{vu}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)$
and
$N(v)=\{X_{v,u}:w_{vu}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)\}$.
Step 1: We color all edges of the subdivision graph of $T$ and certain edges of $C(G)$.
  • We pick $v_{1},v_{2}\in S_{1}(v_{0})$ and assign the pair $(1,1)$ of colors to $X_{v_{0},v_{1}}$ and $X_{v_{0},v_{2}}$. Without loss of generality, we choose $v_{1}\in S_{1}(v_{0})$ such that $v_{1}$ has a neighbor in $S_{2}(v_{0})$. Since the degree of $v_{0}$ is at most $\Delta(G)$ and the pair $(1,1)$ is used twice to color the elements of $N(v_{0})$, we can color every pair of $N(v_{0})\setminus\{X_{v_{0},v_{1}},X_{v_{0},v_{2}}\}$ with different pairs of colors from $(1,1)$. Inspired by an algorithm due to Kalinowski–Pilśniak–Woźniak [9, Proof of Theorem 2.2], we modify the algorithm to color all the edges of the subdivision graph of $T$ so that $v_{0}$ is the only vertex such that the pair $(1,1)$ appears twice in $N(v_{0})$. Hence $v_{0}$ will be fixed by every color-preserving automorphism, and all vertices in $N(v_{0})$ will be fixed, except possibly $v_{1}$ and $v_{2}$. For $i\in\{1,2\}$, let
    $T_{v_{i}}=\{X_{v_{i},u}\in N(v_{i}):u\in S_{2}(v_{0})\}.$
    Note that $T_{v_{1}}\neq\emptyset$ since $v_{1}$ has a neighbor in $S_{2}(v_{0})$, while $T_{v_{2}}$ may possibly be empty.
    In order to distinguish $v_{1}$ and $v_{2}$, we need to take the following steps:
    • ($\ast$)
      We color the elements of $T_{v_{1}}$ and $T_{v_{2}}$ with the same pairs so that none of them contains the pair $(1,1)$.
    • ($\ast\ast$)
      If $X_{v_{1},u}\in T_{v_{1}}$, then $u\in S_{2}(v_{0})$. Hence $v_{0}u\in E(\overline{G})\subseteq E(C(G))$. Let
      $Z_{v_{1}}=\{v_{0}u\in E(C(G)):X_{v_{1},u}\in T_{v_{1}}\}.$
      We color all edges of $Z_{v_{1}}$ with $1$.
      Similarly, we color all edges of the set
      $Z_{v_{2}}=\{v_{0}u\in E(C(G)):X_{v_{2},u}\in T_{v_{2}}\}$
      with 0 (if this set is nonempty). In both cases (i.e., $Z_{v_{2}}=\emptyset$ or $Z_{v_{2}}\neq\emptyset$), the resulting color pattern in $C(G)$ distinguishes $v_{1}$ and $v_{2}$.
Claim 3.5.
If $f$ is an edge coloring of $C(G)$ and $v_{0}$ is fixed by a $f$-preserving automorphism $\phi$, then $\phi(v_{i})=v_{i}$ for $i\in\{1,2\}$.
Proof.
Otherwise, $\phi(v_{1})=v_{2}$. Therefore, if $u\in S_{2}(v_{0})\cap N_{T}(v_{1})$, where $N_{T}(v)$ is the open neighborhood of a vertex $v$ in the tree $T$, then $\phi(u)\in S_{2}(v_{0})\cap N_{T}(v_{2})$. Since $uv_{0}\in E(C(G))$ and $f(uv_{0})=1$, we have $\phi(u)\phi(v_{0})\in E(C(G))$ and $f(\phi(u)\phi(v_{0}))=f(\phi(u)v_{0})=1$. This contradicts $(\ast\ast)$. ∎
  • Suppose that for some integer $k$, every element $u\in\bigcup_{i\leq k}S_{i}(v_{0})$ is colored so that it is fixed by every color-preserving automorphism. For every $u\in S_{k}(v_{0})$, color $X_{u,z}$, where $z$ is a child of $u$ in $T$, with distinct color pairs except $(1,1)$. Hence, all neighbors in $N(u)\cap S_{k+1}(v_{0})$ will be fixed. So, the elements of $\bigcup_{i\leq k+1}S_{i}(v_{0})$ are also fixed by every color-preserving automorphism. Continuing in this fashion, we color all the edges of $S(T)$.
Step 2: We color all the remaining edges in $C(G)$ with 0.
Step 3: Using Lemmas 2.1 and 2.2, we prove that the coloring we defined above (say, $c$) is preserved only by $id_{C(G)}$.
  • Let $\psi\in Aut(C(G))$ be an automorphism that preserves $c$. By Lemma 2.2, $\psi(V_{i})=V_{i}$ for $i\in\{1,2\}$. Since all vertices of $T$ are fixed by $\psi$ and the vertices of $T$ correspond to $V_{1}$, it follows that $\psi(u)=u$ for every $u\in V_{1}$. By Lemma 2.1, we conclude that $\psi=id_{C(G)}$.
Case (iii): $G$ is a tree.
Then $G$ has a central vertex or a central edge that is fixed by every automorphism of $G$.
  1. Case A: Suppose $G$ has a central vertex $v_{0}$. If $S_{2}(v_{0})\neq\emptyset$, we consider $G$ rooted in $v_{0}$ and apply the coloring argument of Case (ii), which fixes vertices successively according to their distance from $v_{0}$. If $S_{2}(v_{0})=\emptyset$, then $G\cong K_{1,n-1}$. In this case, it is known that
    $D^{\prime}(C(K_{1,n-1}))=\left\lceil\sqrt{n-1}\right\rceil=\left\lceil\sqrt{\Delta(K_{1,n-1})}\right\rceil.$
    Thus, the desired bound holds.
  2. Case B: If $G$ has a central edge $e_{0}=xy$, let $T_{1}$, $T_{2}$ be subtrees obtained by deleting $e_{0}$. If we put different colors on the edges $xw_{xy}$ and $w_{xy}y$, then $x$ and $y$ are fixed by every automorphism. We color $T_{i}$ using the same method as in Case A for $i\in\{1,2\}$.
Proposition 3.6.
For all integers $m,n\geq 3$,
$D^{\prime}(C(K_{m,n}))\leq\max(D^{\prime}(K_{m}),D^{\prime}(K_{n}))+1.$
Proof.
Let $V=\{v_{1},\dots,v_{n}\}$ and $U=\{u_{1},\dots,u_{m}\}$ be the bipartitions of $K_{m,n}$, and let $G=C(K_{m,n})$. Then $G[V]\cong K_{n}$ and $G[U]\cong K_{m}$ where $G[V]$ and $G[U]$ are the induced subgraphs of $G$ with respect to $V$ and $U$ respectively. Let $G_{n}:=G[V]$ and $G_{m}:=G[U]$ (see Figure 4).
[…]
Figure 4. A schematic representation of $C(K_{n,m})$. Representative edges of the cliques $K_{n}$ and $K_{m}$ are drawn, and each edge $v_{i}u_{j}$ is subdivided by a vertex $w_{v_{i}u_{j}}$.
Suppose that $f_{n}:E(G_{n})\to\{1,\dots,D^{\prime}(K_{n})\}$ and $f_{m}:E(G_{m})\to\{1,\dots,D^{\prime}(K_{m})\}$ are distinguishing edge colorings of $G_{n}$ and $G_{m}$, respectively. Let $d=\max(D^{\prime}(K_{m}),D^{\prime}(K_{n}))$. Define a coloring $f:E(G)\to\{1,\dots,d+1\}$ as follows:
[…]
Claim 3.7.
$f$ is a distinguishing edge coloring of $G$.
Proof.
Let $\phi\in\mathrm{Aut}(G)$ be an automorphism that preserves $f$. By Lemma 2.2, $\phi(V\cup U)=V\cup U$ and $\phi(V_{2})=V_{2}$. For each $v_{i}\in V$, all edges of the form $v_{i}w_{v_{i}u_{j}}$ are colored $d+1$ with respect to $f$, whereas for each $u_{j}\in U$, no edge of the form $w_{v_{i}u_{j}}u_{j}$ is colored $d+1$. Thus, the property
[…]
characterizes vertices of $V$. Since $\phi$ preserves $f$, it follows that $\phi(V)=V$ and $\phi(U)=U$. Define $\psi_{V}:=\phi\restriction V$ and $\psi_{U}:=\phi\restriction U$, i.e., the restrictions of $\phi$ to $V$ and $U$ respectively. It is clear that $\psi_{V}\in Aut(G_{n})$ and $\psi_{U}\in Aut(G_{m})$. [^2] Let $v_{i}v_{j}\in E(G_{n})$. Then $v_{i}v_{j}\in E(G)$ and both endpoints $v_{i}$ and $v_{j}$ lie in $V$. Since $\phi\in Aut(G)$, we have $\phi(v_{i})\phi(v_{j})\in E(G)$. Since $\phi(V)=V$, it follows that $\phi(v_{i}),\phi(v_{j})\in V$, and hence $\phi(v_{i})\phi(v_{j})\in E(G_{n})$. Thus, $\psi_{V}$ preserves adjacency in $G_{n}$, and hence $\psi_{V}\in Aut(G_{n})$. Similarly, $\psi_{U}\in Aut(G_{m})$.
Fix any $v_{i}v_{j}\in E(G_{n})$. Then $\psi_{V}(v_{i})\psi_{V}(v_{j})\in E(G_{n})$ where $\psi_{V}(v_{i})$ and $\psi_{V}(v_{j})$ belong to $V$. Since $\psi_{V}\in Aut(G_{n})$ and automorphisms preserve adjacency, the endpoints of $\psi_{V}(v_{i}v_{j})$ are $\psi_{V}(v_{i})$ and $\psi_{V}(v_{j})$. Now $G_{n}\cong K_{n}$ is a simple graph. Thus, we have
$\psi_{V}(v_{i}v_{j})=\psi_{V}(v_{i})\psi_{V}(v_{j}).$
Since $\phi$ preserves $f$, we obtain
$f_{n}(\psi_{V}(v_{i}v_{j}))=f_{n}(\psi_{V}(v_{i})\psi_{V}(v_{j}))=f(\psi_{V}(v_{i}v_{j}))=f(\phi(v_{i}v_{j}))=f(v_{i}v_{j})=f_{n}(v_{i}v_{j}).$
Thus, $\psi_{V}$ preserves $f_{n}$. Similarly, $\psi_{U}$ preserves $f_{m}$. We know $f_{n}$ and $f_{m}$ are distinguishing edge colorings of $G_{n}$ and $G_{m}$ respectively. Thus, $\psi_{V}$ and $\psi_{U}$ fixes all vertices of $V$ and $U$ respectively. Consequently, $\phi$ fixes $V\cup U$ pointwise. By Lemma 2.1, we conclude that $\phi=id_{C(K_{m,n})}$. ∎
Theorem 3.8.
Let $G$ be a connected graph of order $n\geq 4$. Then $D(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$.
Proof.
In view of Fact 3.1(1), it suffices to show that $D(C(G))\leq D^{\prime\prime}(G)$. Let $f:V(G)\cup E(G)\to\{1,...,D^{\prime\prime}(G)\}$ be a total coloring of $G$ (not necessarily proper) that is preserved only by $id_{G}$. We define $\tilde{f}:V(C(G))\to\{1,...,D^{\prime\prime}(G)\}$ so that $\tilde{f}(x)=f(x)$ for all $x\in V_{1}$ and $\tilde{f}(w_{xy})=f(xy)$ for all $w_{xy}\in V_{2}$. We claim that $\tilde{f}$ is preserved only by $id_{C(G)}$. Let $\psi$ be a $\tilde{f}-$preserving automorphism of $C(G)$. If $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$ where $\pi_{\alpha}$ is the induced automorphism of $\alpha\in Aut(G)$, which is defined in the proof of Lemma 2.3, then $H=Aut(C(G))\cong Aut(G)$ (see the proof of Lemma 2.3). Thus, $\psi=\pi_{\alpha}$ for some $\alpha\in Aut(G)$ where $\psi(V_{1})=V_{1}$ and $\psi(V_{2})=V_{2}$. Since $\psi$ is a $\tilde{f}-$preserving automorphism of $C(G)$, we can see that $\alpha$ preserves $f$ for the following reasons:
  1. (1)
    If $v\in V_{1}$, then
    $\tilde{f}(\psi(v))$ $=\tilde{f}(v)$
    $\implies f(\psi(v))=f(v)$
    $\implies f(\alpha(v))=f(\pi_{\alpha}(v))=f(v).$
  2. (2)
    If $w_{xy}\in V_{2}$, then
    $\tilde{f}(\psi(w_{xy}))$ $=\tilde{f}(w_{xy})$
    $\implies\tilde{f}(\pi_{\alpha}(w_{xy}))=f(xy)$
    $\implies\tilde{f}(w_{\alpha(x)\alpha(y)})=f(xy)$
    $\implies f(\alpha(x)\alpha(y))=f(xy)$
    $\implies f(\alpha(xy))=f(xy).$
So, $\alpha=id_{G}$ since $f$ is preserved only by $id_{G}$. Consequently, $\psi=\pi_{\alpha}=id_{C(G)}$. ∎
[…]
Figure 5. A vertex coloring of $C(K_{1,4})$ with $2$ colors and an edge coloring of $C(K_{1,5})$ with $3$ colors that are only preserved by trivial automorphisms.
Remark 3.9.
For any integer $n\geq 1$, we have
$D^{\prime}(C(K_{1,n}))=D(C(K_{1,n}))=\lceil\sqrt{n}\rceil=\lceil\sqrt{\Delta(K_{1,n})}\rceil.$
Thus, there exist graphs for which the bounds mentioned in Theorems 3.4 and 3.8 cannot be improved (see Fig. 5).
Theorem 3.10.
Let $G$ be a connected graph of order $n$. The following holds:
  1. (1)
    If $n\geq 3$, then $D(M(G))\leq\Delta(G)$. Moreover, the bound is sharp.
  2. (2)
    $D^{\prime}(M(G))\leq 3$ for any positive integer $n$.
Proof.
Let $E$ be the set of all endline edges of $G^{+}$.
(1). By Facts 3.1(4) and 2.5 (1), $D(M(G))=D(L(G^{+}))=D^{\prime}(G^{+})$.
Case(i): $G\not\cong C_{n}$ for any $n\geq 3$. We show that $D^{\prime}(G^{+})\leq D^{\prime}(G)$. Then $D^{\prime}(G^{+})\leq\Delta(G)$ by Fact 3.1(5). Let $g:E(G)\rightarrow\{1,...,D^{\prime}(G)\}$ be a coloring of $E(G)$ that is preserved only by $id_{G}$. Define the mapping $g^{\prime}:E(G^{+})\rightarrow\{1,...,D^{\prime}(G)\}$ as follows:
$g^{\prime}(e)=\begin{cases}1&\text{if}\,e\in E,\\ g(e)&\text{if}\,e\not\in E.\end{cases}$
We show that $g^{\prime}$ is preserved only by $id_{G^{+}}$. If $\psi\in Aut(G^{+})$ preserves $g^{\prime}$, then $\psi$ maps edges from $E$ to $E$ by Lemma 2.7. So, $\psi\restriction G\in Aut(G)$ preserves $g$. Thus, $\psi\restriction G=id_{G}$. By Lemma 2.6, $\psi=id_{G^{+}}$.
Case(ii): $G\cong C_{n}$ for some $n\geq 3$.
Claim 3.11.
$D^{\prime}(G^{+})=2$.
Proof.
Let $U_{1}=V(G)=\{v_{1},\ldots,v_{n}\}$ and $U_{2}=V(G^{+})\setminus V(G)=\{u_{1},\ldots,u_{n}\}$, where $u_{i}$ is the pendant vertex adjacent to $v_{i}$. Since the graph $G^{+}$ admits a nontrivial automorphism, any edge coloring with one color is preserved by a non-identity automorphism. Hence $D^{\prime}(G^{+})\geq 2$. Thus, it is enough to show that $D^{\prime}(G^{+})\leq 2$. Define a coloring $c:E(G^{+})\to\{1,2\}$ as follows:
$c(e)=\begin{cases}2&\text{if }e=v_{1}v_{2}\text{ or }e=v_{1}u_{1},\\ 1&\text{otherwise}.\end{cases}$
We show that $c$ is a distinguishing edge coloring.
  • Let $\psi\in\mathrm{Aut}(G^{+})$ be a $c$-preserving automorphism. By Lemma 2.7, $\psi(U_{i})=U_{i}$ for $i=1,2$. Hence $\psi$ maps $E(G)$ to $E(G)$ and the set of endline edges $E$ to itself.
  • The edge $v_{1}u_{1}$ is the unique endline edge colored $2$. Therefore $\psi(v_{1}u_{1})=v_{1}u_{1}$. Since $u_{1}$ is the only vertex of degree $1$ incident with this edge, it follows that $\psi(u_{1})=u_{1}$ and consequently $\psi(v_{1})=v_{1}$.
  • The edge $v_{1}v_{2}$ is colored differently from any other edge in the cycle graph $G$. Thus $\psi(v_{1}v_{2})=v_{1}v_{2}$. Since $\psi(v_{1})=v_{1}$, it follows that $\psi(v_{2})=v_{2}$.
  • For $v_{p}\in V(G)\setminus\{v_{1},v_{2}\}$, let $P_{v_{1},v_{2},v_{p}}$ denote the unique path in the cycle between $v_{1}$ and $v_{p}$ that contains $v_{2}$. Now, $\psi$ fixes all the vertices of $G$ since for any $v_{p},v_{q}\in V(G)\backslash\{v_{1},v_{2}\}$ such that $v_{p}\neq v_{q}$, the length of $P_{v_{1},v_{2},v_{p}}$ is different from the length of $P_{v_{1},v_{2},v_{q}}$.
  • Thus, $\psi(u)=u$ for every $u\in U_{1}$. By Lemma 2.6, we conclude that $\psi=id_{G^{+}}$.
Hence, $D^{\prime}(G^{+})\leq 2$. This concludes the proof of the claim. ∎
Thus, $D(M(G))=D^{\prime}(G^{+})=2=\Delta(G)$. This also shows that the bound mentioned in (1) is sharp.
(2). Since every line graph is claw-free, and the distinguishing index of a connected claw-free graph is at most $3$ (see [16, Theorem 3.4]), we have $D^{\prime}(M(G))=D^{\prime}(L(G^{+}))\leq 3$. ∎
Data availability No data are associated with this article.
Declarations
Conflict of interest The author declares that there is no conflict of interest.